# CAML: Collaborative Auxiliary Modality Learning for Multi-Agent Systems

### v1

@@ 3 Collaborative Auxiliary Modality Learning
Specifically, in the teacher model, each agent has access to all multimodal observations and independently processes its local observations to produce embeddings. These embeddings are then shared among agents based on whether the system operates in a centralized or decentralized manner. If the system is centralized, all collaborative agents share their embeddings with one designated ego agent for centralized processing. If the system is decentralized, each agent shares the embeddings with other agents. We provide a detailed complexity analysis of CAML for both operation manners in the Appendix A.4.2. Subsequently, the shared embeddings corresponding to the same modality are aggregated together. Then we fuse (e.g., via concatenation or cross-attention) the aggregated embeddings of different modalities to create a comprehensive multimodal embedding. This multimodal embedding is then passed through a prediction module to produce the teacher model’s final prediction. The student model follows a similar network architecture as the teacher. However, instead of processing all modalities, each agent processes only a single or a subset of modalities, which can vary across agents. By sharing these embeddings among agents, the student model also constructs a multimodal embedding, leveraging the different modalities observed by various agents. This multimodal embedding is then used to generate the student model’s prediction. Thus, our approach enables the student model to maintain strong predictive performance despite missing modalities during testing, significantly enhancing its robustness and generalizability.
Figure 2: CAML Approach Pipeline. The teacher model (top) aggregates and shares multimodal embeddings across agents for prediction. In contrast, the student model (bottom) processes a subset of modalities per agent and shares them to form a multimodal embedding. This allows the student model to handle missing modalities during testing, while still generating robust predictions. Please see details in Section 3.
4 Analysis
To compare whether CAML outperforms AML with a single agent theoretically, we analyze from two key perspectives: uncertainty reduction and data coverage enhancement. Data coverage can be further discussed from two dimensions: complementary information and information gain. We aim to address three major questions: (a) Uncertainty Reduction: Does the collaboration among multiple agents help reduce the variance of the posterior distribution, resulting in more confident estimates? (b) Complementary Information: Does the collaboration of multiple agents provide complementary information that increases data coverage? Specifically, does combining observations from each agent lead to a more accurate and comprehensive prediction compared to using a single agent? (c) Information Gain: Does the collaboration increase the mutual information between the observations and the true label?
Uncertainty Reduction.
To address question (a) about uncertainty reduction, the prior $P(y)$ is typically assumed to be Gaussian: $P(y)={\mathcal{N}}(y|\mu_{0},\sigma_{0}^{2})$, where $\mu_{0}$ is the prior mean and $\sigma_{0}^{2}$ is the prior variance.
Single-Agent. In the single-agent case, we assume that only agent ${\mathcal{A}}_{i}$ is available and its likelihood $P(x_{i}|y)$ is Gaussian: $P(x_{i}|y)={\mathcal{N}}(x_{i}|\mu_{i}(y),\sigma_{i}^{2})$, where $\mu_{i}(y)$ is the mean of the observation $x_{i}$ given $y$, $\sigma_{i}^{2}$ is the variance of the agent ${\mathcal{A}}_{i}$’s observations. The posterior distribution $P(y|x_{i})$ is proportional to the product of the prior and likelihood, $P(y|x_{i})\propto P(y)P(x_{i}|y)$, which is also Gaussian. And the posterior variance $\sigma_{single}^{2}=\big{(}\frac{1}{\sigma_{0}^{2}}+\frac{1}{\sigma_{i}^{2}}\big{)}^{-1}$ (Murphy, 2007).
Multi-Agent. In the case of multi-agent collaboration, we model the joint likelihood of the observations $X$ as a multivariate Gaussian distribution, conditioned on the true target variable $y$: $P(X|y)={\mathcal{N}}(X|\mu_{X}(y),\Sigma)$, where $\mu_{X}(y)$ is the joint mean of the observations from all agents, conditioned on $y$, $\Sigma$ is the covariance matrix, encoding the correlations between the observations from multiple agents. The posterior $P(y|X)\propto P(y)P(X|y)$, is another Gaussian, with variance $\sigma_{multi}^{2}=\big{(}\frac{1}{\sigma_{0}^{2}}+\bf 1^{T}\Sigma^{-1}\bf 1\big{)}^{-1}$ (Murphy, 2007). Since $\bf 1^{T}\Sigma^{-1}\bf 1\geq\frac{1}{\sigma_{i}^{2}}$ for any $i$, we have $\sigma^{2}_{multi}\leq\sigma_{single}^{2}$. In the extreme case where all agents’ observations are perfectly correlated (e.g., they all observe the same thing), the posterior variance would be equivalent to that of a single agent. However, multi-agent collaboration reduces variance compared to a single agent, as long as the observations are not perfectly correlated, proving that collaboration reduces uncertainty.
Enhanced Data Coverage.
In comparing data coverage between CAML and AML, we analyze it from two key aspects: complementary information and information gain.
Complementary Information.
To address question (b) about complementary information, we study data coverage and information provided by each agent in a multi-agent system. Let the entire data space be denoted as ${\mathcal{D}}$, which consists of various subsets. Each agent ${\mathcal{A}}_{i}$ in the system covers a subset of this data space: ${\mathcal{C}}_{i}\subseteq{\mathcal{D}}$. The overall coverage by the system is given by the union of all subsets covered by individual agents: ${\mathcal{C}}_{multi}=\cup_{i=1}^{N}{\mathcal{C}}_{i}$. This ensures that $|{\mathcal{C}}_{multi}|\geq\max|{\mathcal{C}}_{i}|$. If only a single agent is available, it can only observe a portion of the data space, leaving parts of the space unobserved, which leads to incomplete information for estimating the true label $y$. We show an qualitative example of multi-agent collaboration provides complementary information to enhance data coverage in Figure 7 in the Appendix.
From a probabilistic perspective, when multi-agent collaboration is in place, the combined likelihood $P(X|y)$ is modeled as a multivariate distribution (as discussed in Section 4). This approach provides a broader and more accurate representation of the data space by integrating information from all agents and modeling the dependencies and correlations between them. Compared to a univariate distribution $P(x_{i}|y)$ for a single agent ${\mathcal{A}}_{i}$, the multivariate distribution covers a larger portion of the data space ${\mathcal{D}}$, thus enhancing data coverage. This allows the exploration of more complex patterns, relationships, and complementary information from different agents. By capturing a richer set of interactions and correlations among the agents’ observations, the multivariate distribution supports more informed decision-making. The model’s predictions are based on a comprehensive view of the environment, thus leading to more accurate outcomes.
Information Gain.
To address question (c) about information gain, we analyze using information theory. Let $I(y;x_{i})$ represent the mutual information between the true label $y$ and agent ${\mathcal{A}}_{i}$’s observation $x_{i}$, which quantifies how much information $x_{i}$ provides about the estimation of $y$. The mutual information between $y$ and the set of all observations $X$ is $I(y;X)$. In the context of multi-agent collaboration, the joint observations $X$ from multiple agents typically provide more comprehensive information about the true label $y$ compared to the observation of any single agent. Therefore, the mutual information $I(y;X)$ is always greater than or equal to the mutual information from a single agent: $I(y;X)\geq I(y;x_{i})$. Thus, the combined observations from multi-agent collaboration provide more information about $y$ than a single observation, improving the overall estimate. By leveraging the combined knowledge from multiple agents, the prediction of $y$ becomes more accurate, reflecting added value of collaboration. Multi-agent systems are generally more informative, as the interaction and joint information between agents can reduce uncertainty about the target variable, as discussed in Section 4.
@@ 5.1 Collaborative Decision-Making
Following prior research (Cui et al., 2022; Gao et al., 2024), we focus on three complex traffic scenarios prone to accidents due to limited sensor coverage or obstructed views, as illustrated in Fig. 6. For more details about the scenarios, please refer to the Appendix A.1. For each scenario, we collect 24 trials, dividing them into 12 trials for training through behavior cloning (BC) and 12 trials for testing. Each trial includes RGB images and LiDAR point clouds captured by a variable number of connected vehicles, along with the ground truth actions of the ego vehicle. At each timestamp, the ego vehicle has a maximum of three collaborative vehicles, provided their distance is within a threshold of 150 meters (Gao et al., 2024; Cui et al., 2022). For each vehicle, both RGB and LiDAR data are used during training, while only RGB data is used during testing in CAML.
@@
For processing RGB data, we first resize the image to $224\times 224$ and use ResNet-18 (He et al., 2016) as the encoder to extract a feature map. We then apply self-attention on the feature map to dynamically compute the importance of features at different locations. After the self-attention, we apply three convolution layers with each followed by a ReLU activation. Finally, we obtain a 256-dimensional feature representation after passing through a fully connected layer. To facilitate the collaboration and aggregation of RGB feature embeddings from connected vehicles to the ego vehicle, we use the cross-attention mechanism. For processing the LiDAR data, we use the Point Transformer (Zhao et al., 2021) as the encoder and utilize the COOPERNAUT (Cui et al., 2022) model to aggregate LiDAR feature embeddings.
For the training of Knowledge Distillation (KD), we first train a teacher model offline using a binary cross-entropy loss, where each vehicle has both RGB and LiDAR data. Then we train a student model to mimic the behavior of the teacher model with only RGB data for each vehicle. For each data point, the student model receives the same RGB image that the teacher model was given. For further details on the KD training process, please refer to Appendix A.5. For the prediction module, we use a three-layer MLP. And for the detailed training settings, please see Appendix A.4.1. We employ the following two metrics for evaluation: (1) Accident Detection Rate (ADR): This is the ratio of accident-prone cases correctly detected by the model compared to the total ground truth accident-prone cases. An accident-prone case is identified when the ego vehicle performs a braking action. This metric measures the model’s effectiveness in identifying potential accidents. (2) Expert Imitation Rate (EIR): This denotes the percentage of actions accurately replicated by the model out of the total expert actions. It serves to evaluate how well the model mimics expert driving behavior.
@@
We implement the following baselines for comparison: (1) AML (Shen et al., 2023): In the AML setting, the ego vehicle operates independently without collaboration with other vehicles (non-collaborative). Both RGB and LiDAR data are available during training for the vehicle, while only RGB data is available during testing. (2) COOPERNAUT (Cui et al., 2022) (Single-Agent): Processes LiDAR data during both training and testing. COOPERNAUT uses the Point Transformer (Zhao et al., 2021) as the backbone, encoding raw 3D point clouds into keypoints. (3) STGN (Gao et al., 2024) (Single-Agent): Utilizes spatial-temporal graph networks for decision-making, with RGBD data used for both training and testing.
@@
Figure 3: Performance Comparison of CAML Against Baselines. We evaluate performance using two metrics: Accident Detection Rate (ADR) and Expert Imitation Rate (EIR) across three accident-prone scenarios: (a) Overtaking, (b) Left Turn, and (c) Red Light Violation. The baselines, AML, COOPERNAUT, and STGN, operate in a single-vehicle, non-collaborative setting. In contrast, CAML demonstrates superior performance across all scenarios compared to these baselines by up to 58.13%, benefiting considerably from the multi-agent collaboration.
Baselines Comparison.
How well does CAML perform against other methods for decision-making in CAV? We evaluate CAML against the baselines and present the results in Figure 3, which demonstrate a clear performance advantage of CAML across all three accident-prone scenarios. The evaluation metrics, accident detection rate (ADR) and expert imitation rate (EIR), reveal that CAML consistently outperforms AML, COOPERNAUT, and STGN. In particular, CAML achieves notable improvements in ADR compared to AML: ${\bf 13.2\%}$ in the overtaking scenario, ${\bf 32.6\%}$ in the left turn scenario, and a significant ${\bf 58.13\%}$ in the red light violation scenario. The more pronounced improvements in the left turn and red light violation scenarios can be attributed to the higher complexity of these situations, where restricted views and occlusions present greater challenges for decision-making. Unlike the overtaking scenario on a two-way road, which is relatively less constrained, left turns and red light violations often involve more unpredictable vehicle and pedestrian interactions, requiring enhanced situational awareness. In these more demanding cases, the collaborative framework of CAML proves especially beneficial, as it allows the ego vehicle to aggregate additional sensory data from connected vehicles, significantly boosting its capacity to detect potential accidents and respond proactively, such as applying braking when necessary to avoid collisions.
As detailed in Section 4, the collaborative nature of CAML plays a critical role in reducing the uncertainty in the decision-making processes. By incorporating sensory data from multiple connected vehicles, CAML can draw on a richer and more diverse dataset, which enables more reliable predictions. This collaborative approach not only reduces the uncertainty in estimations but also enhances data coverage by leveraging complementary information from all connected agents. As a result, the ego vehicle is able to form a more accurate and comprehensive understanding of its environment, particularly in scenarios where its own sensing capabilities are limited by obstructions or blind spots. Compared to single-agent systems, where decisions rely solely on local sensory data, the multi-agent collaboration in CAML allows the ego vehicle to better handle complex driving environments, especially in accident-prone situations. These baseline comparison results of improvements in safety and decision-making align well with our theoretical analysis.
@@
Figure 4: Modality-Efficient Superiority of CAML Against STGN with Multi-Agent Settings. We compare CAML with STGN with multi-agent settings, using ADR and EIR metrics across three accident-prone scenarios: (a) Overtaking, (b) Left Turn, and (c) Red Light Violation. While STGN uses both RGB and depth data during testing, CAML relies solely on RGB, yet achieves comparable, or even better performance. This highlights the effectiveness of CAML, leveraging LiDAR as an auxiliary modality during training to enhance performance.
(a) Overtaking
(b) Left Turn
(c) Red Light Violation
@@
Figure 5: System Generalizability of CAML. We evaluate the generalizability of CAML by testing the case where we have multi-agent collaboration during training, but only a single agent during testing. The performance is assessed using ADR and EIR across three accident-prone scenarios: (a) Overtaking, (b) Left Turn, and (c) Red Light Violation. CAML with a single agent during testing consistently outperforms the three baselines across all scenarios, offering a valuable and cost-effective solution for practical applications.
Modality-Efficient Superiority.
How does CAML compare with other approaches that have access to more modalities during testing? By modality-efficient superiority, we refer to a model’s ability to achieve comparable or even superior performance using fewer modalities compared to other approaches that rely on a richer set of modalities. We evaluate CAML against STGN (Gao et al., 2024) with multi-agent settings. CAML uses only RGB data during testing but STGN uses both RGB and depth data. Both models are evaluated using the same metrics, ADR and EIR, across the three accident-prone scenarios. Despite the fact that STGN utilizes both RGB and depth data during testing, CAML achieves comparable, and in some cases superior, performance while relying solely on RGB data, as illustrated in Figure 4. Notably, CAML exceeds the ADR of STGN by ${\bf 9.26\%}$ in the left-turn scenario, demonstrating that our model can enhance driving safety even when constrained to fewer modalities. This further underscores the strength of CAML, which effectively leverages LiDAR data as an auxiliary modality during training to boost performance, even when such data is unavailable during testing. The fact that CAML matches or exceeds the performance of a model that uses more data at test time highlights the efficacy of our multi-agent collaboration approach.
System Generalizability.
How effectively does the system generalize when we have fewer agents during testing compared to training? (e.g., we have multi-agent collaboration during training but only single agent during testing). We test the case where multi-agent collaboration is used during training, but only a single agent is present during testing. This test is also motivated by practical constraints, where in many real-world situations, multi-vehicle connected systems are not available, we only have a single vehicle. But it is reasonable to have multi-vehicle connected systems with multiple modalities during training to develop robust models. After training, we can then apply the model on a single vehicle for inference or testing, which is very valuable in practice and provides a cost-effective solution.
We compare the performance with other baselines, using the same evaluation metrics of ADR and EIR, across three accident-prone scenarios. The comparison results are presented in Figure 5. CAML with a single agent during testing outperforms the three baselines across all scenarios, for both ADR and EIR metrics. This demonstrates that even with single agent during testing, CAML remains highly effective, by utilizing the multi-agent collaboration and auxiliary modalities provided by the teacher model during training.
@@
Overall, the experimental results clearly illustrate the superiority of our CAML framework. The ability of CAML to learn a more effective driving policy stems from the collaborative behavior of multiple agents, which together capture a wider and more nuanced representation of data. This broader data coverage enables the ego vehicle to make better-informed decisions, improving safety and performance, particularly in complex, dynamic, and accident-prone environments where isolated agents with limited sensing.

### v2

@@ 3 Collaborative Auxiliary Modality Learning
Specifically, in the teacher model, each agent has access to all multi-modal observations and independently processes its local observations to produce embeddings. These embeddings are then shared among agents based on whether the system operates in a centralized or decentralized manner. If the system is centralized, all collaborative agents share their embeddings with one designated ego agent for centralized processing. If the system is decentralized, each agent shares the embeddings with nearby agents. Subsequently, the shared embeddings corresponding to the same modality are aggregated together. Then we fuse (e.g., via concatenation or cross-attention) the aggregated embeddings of different modalities to create a comprehensive multi-modal embedding. This multi-modal embedding is then passed through a prediction module to produce the teacher model’s final prediction. The student model follows a similar network architecture as the teacher. However, instead of processing all modalities, each agent processes only a single or a subset of modalities, which can vary across agents. By sharing these embeddings among agents, the student model also constructs a multi-modal embedding, leveraging the different modalities observed by various agents. This multi-modal embedding is then used to generate the student model’s prediction. Our approach enables the student model to maintain effective performance despite missing modalities during testing, significantly enhancing its robustness.
We further provide an intuitive analysis offering insights and explanations to justify the effectiveness of CAML by examining how our approach enhances data coverage and leads to greater information gain compared to single-agent approaches. Specifically, we show that aggregating observations from multiple agents enables broader coverage of the data space, capturing complementary information that would be missed by an individual agent. From an information theory perspective, we demonstrate that joint observations from multiple agents generally yield higher mutual information with respect to the ground truth, providing a more informative signal. Please refer to Appendix A.1 for more details.
System Complexity.
After designing the CAML framework, we present a detailed complexity analysis. If the system is centralized, all collaborative agents share their data with one designated ego agent for centralized processing. Each of the $N-1$ collaborative agents performs its local computation independently, with a time complexity of $O(T_{c})$ and a space complexity of $O(S_{c})$, where $T_{c}$ represents the time required for local computation, and $S_{c}$ is the associated space. Thus, the total computation time and space complexities for all collaborative agents are $O(T_{c}(N-1))$ and $O(S_{c}(N-1))$, respectively. For simplicity, assuming each communication from one collaborative agent to the ego agent consumes $O(D)$ time complexity and $O(M)$ space complexity, where $D$ is the time required for communication and $M$ is the corresponding space. Therefore, the total communication time and space complexities for gathering information at the ego agent are $O(D(N-1))$ and $O(M(N-1))$, respectively. Then the ego agent aggregates the received data, running a model, having a time and space complexity $O(T_{e})$ and $O(S_{e})$, where $T_{e}$ and $S_{e}$ represent the time and space required for the ego agent’s computation. So the total time and space complexities are $O(T_{c}(N-1)+D(N-1)+T_{e})$ and $O(S_{c}(N-1)+M(N-1)+S_{e})$, respectively.
If the system is decentralized, each agent performs its local computation and shares information with nearby agents. For simplicity, let the local computation for a single agent has a time complexity of $O(T)$, where $T$ is the time required for local computation. Assume that communication from one agent to another agent requires $O(D)$ time complexity and $O(M)$ space complexity, where $D$ represents the time of communication between two agents, and $M$ denotes the space required for such communication. For $N$ agents, the total computation time complexity is $O(NT)$. In the worst case, each agent share data with all other agents, this can result in $O(N^{2}D)$ for pairwise sharing. So the total time complexity is $O(NT+N^{2}D)$. For space complexity, the storage requirement for all agents is $O(NS)$, where $S$ is the space needed per agent. Communication between agents adds an additional complexity of $O(N^{2}M)$. So the total space complexity is $O(NS+N^{2}M)$. In the typical case, if each agent communicates with only other $k$ agents ($k\ll N$) rather than all $N-1$ agents. The total time and space complexities become $O(NT+NkD)$ and $O(NS+NkM)$, respectively.
@@ 4.1 Collaborative Decision-Making
The dataset is generated using the AUTOCASTSIM benchmark [29], which features three complex and accident-prone traffic scenarios designed for CAV. These scenarios are characterized by limited sensor coverage and obstructed views, as illustrated in Fig. 6. Following prior works COOPERNAUT [5] and STGN [8], we employ an expert agent with full access to the traffic scene information to collect data. This expert agent is deployed in the environment to generate the data trails, where each trail captures RGB images and LiDAR point clouds from both the ego vehicle and collaborative vehicles, along with the control actions of the ego vehicle. In line with the setup used in STGN [8], the ego vehicle can have up to three collaborative vehicles at each timestep, provided they are within a 150-meter radius. Each trail represents a unique instantiation of the traffic scenario, with differences arising from the randomized scenario configurations. For further dataset details, please refer to Appendix A.2.1.
@@
After data collection, we train the models using Behavior Cloning (BC), following the setup established by prior works COOPERNAUT [5] and STGN [8]. For each vehicle, both RGB and LiDAR data are used during training, while only RGB data is used during testing in CAML. Please refer to Appendix A.2.3 for more details on the BC procedure. All experiments are conducted across four repeated runs, and we report the mean and standard deviation of the results.
For processing RGB data, we first resize the image to $224\times 224$ and use ResNet-18 [14] as the encoder to extract a feature map for each vehicle. We then apply self-attention on the feature map to dynamically compute the importance of features at different locations. After the self-attention, we apply three convolution layers with each followed by a ReLU activation. Finally, we obtain a 256-d feature representation after passing through a fully connected layer. To aggregate RGB feature embeddings from connected vehicles to the ego vehicle, we use the cross-attention mechanism. For processing the LiDAR data, we use the Point Transformer as the encoder for each vehicle and utilize the COOPERNAUT [5] model to aggregate LiDAR feature embeddings between connected vehicles. Then we concatenate the final RGB and LiDAR embeddings for the ego vehicle’s decision-making, with a three-layer MLP as the prediction module to output the action.
For the training of Knowledge Distillation (KD), we first train a teacher model offline using a binary cross-entropy loss, where each vehicle has both RGB and LiDAR data. Then we train a student model to mimic the behavior of the teacher model with only RGB data for each vehicle. For each data sample, the student model receives the same RGB image that the teacher model was given. For further details on the KD training process, please refer to Appendix A.5. And for the detailed training settings, please see Appendix A.4. We employ the following two metrics for evaluation: (1) Accident Detection Rate (ADR): This is the ratio of accident-prone cases correctly detected by the model compared to the total ground truth accident-prone cases. An accident-prone case is identified when the ego vehicle performs a braking action. This metric measures the model’s effectiveness in identifying potential accidents. (2) Expert Imitation Rate (EIR): This denotes the percentage of actions accurately replicated by the model out of the total expert actions. It serves to evaluate how well the model mimics expert driving behavior.
@@
We implement the following baselines for comparison: (1) AML [32]: In the AML setting, the ego vehicle operates independently without collaboration with other vehicles (non-collaborative). Both RGB and LiDAR data are available during training for the vehicle, while only RGB data is available during testing. (2) COOPERNAUT [5]: A collaborative method that processes LiDAR data during both training and testing. It employs the Point Transformer [39] as the backbone, encoding raw 3D point clouds into keypoints. (3) STGN [8]: A collaborative approach that utilizes spatial-temporal graph networks for decision-making, with RGBD data used for both training and testing.
Baselines Comparison.
How well does CAML perform against other methods for decision-making in CAV? We evaluate CAML against the baselines and present the results in Fig. 3, which demonstrate a clear performance advantage of CAML across all three accident-prone scenarios.
Compared to single-agent systems like AML, CAML achieves notable improvements in mean ADR: ${\bf 13.0\%}$ in the overtaking scenario, ${\bf 34.6\%}$ in the left turn scenario, and a significant ${\bf 58.1\%}$ in the red light violation scenario. These improvements highlight the advantages of CAML’s collaborative framework, which enables the ego vehicle to aggregate complementary sensory data from connected vehicles. This richer, multi-agent perspective significantly enhances situational awareness, allowing the ego vehicle to detect potential accidents and respond proactively, such as braking in time to avoid collisions, especially in scenarios involving occlusions or restricted views. When compared to COOPERNAUT, CAML achieves improvements in mean ADR by up to ${\bf 21.5\%}$, further demonstrating the strength of its collaborative and multi-modal approach.
@@
Figure 3: Performance Comparison of CAML Against Baselines. We evaluate performance using two metrics: Accident Detection Rate (ADR) and Expert Imitation Rate (EIR) across three accident-prone scenarios: (a) Overtaking, (b) Left Turn, and (c) Red Light Violation. CAML demonstrates superior performance across all scenarios compared to these baselines by up to ${\bf 58.1\%}$, benefiting considerably from the multi-modal multi-agent collaboration.
Additionally, how does CAML compare with other approaches that have access to more modalities during testing? We evaluate CAML against STGN [8] under multi-agent settings. CAML uses only RGB data during testing but STGN uses both RGB and depth data. Despite with fewer modality, CAML achieves comparable, and in some cases superior performance while relying solely on RGB data, as illustrated in Fig. 4. Notably, CAML exceeds the mean ADR of STGN by ${\bf 10.9\%}$ in the left-turn scenario, demonstrating that our model can enhance driving safety even when constrained to fewer modalities. This further underscores the strength of CAML, which effectively leverages LiDAR data as an auxiliary modality during training to boost performance. The fact that CAML matches or exceeds the performance of a model that uses more data at test time highlights the modality efficacy of our approach.
System Generalizability.
How effectively does the system generalize when we have fewer agents during testing compared to training? (e.g., we have multi-agent collaboration during training but only single agent during testing). We test the case where multi-agent collaboration is used during training, but only a single agent is present during testing. This test is also motivated by practical constraints, where in many real-world situations, multi-vehicle connected systems are not available, we only have a single vehicle. But it is reasonable to have multi-vehicle connected systems with multiple modalities during training to develop robust models. After training, we can then apply the model on a single vehicle for testing, which is very valuable in practice and provides a cost-effective solution.
We compare the performance of our approach with COOPERNAUT [5] and STGN [8] under single-agent settings, with results shown in Fig. 4. CAML outperforms the other baselines across all scenarios, for both ADR and EIR metrics. This demonstrates that even with a single agent during testing, CAML remains highly effective, by utilizing the multi-agent collaboration and auxiliary modalities provided by the teacher model during training.
Table 1: Communication and inference efficiency comparison. CAML can operate more efficiently with lower communication overhead and faster inference than both [5] and [8].
Approach PS (KB) Latency (ms)
COOPERNAUT [5] 65.5 90.0
STGN [8] 4.9 18.5
CAML 1.0 3.7
Efficiency Analysis.
We evaluate the communication and inference efficiency of CAML compared to other cooperative approaches, COOPERNAUT [5] and STGN [8]. To assess communication bandwidth, we follow STGN [8] and use the shared package size (PS) metric, which measures the size of data exchanged between connected vehicles, serving as an indicator of communication efficiency in collaborative decision-making. For inference efficiency, we measure latency per batch, which reflects the average time required to process a batch during inference. The evaluation results are summarized in Table 1. As shown, CAML achieves lower PS and latency compared to other approaches, highlighting its superior communication and inference efficiency.
@@
Figure 4: Single-Agent System Generalizability of CAML. We evaluate the generalizability of CAML by testing the case where we have multi-agent collaboration during training, but only a single agent during testing. We compare the performance of our approach with COOPERNAUT and STGN under single-agent settings. CAML with a single agent during testing consistently outperforms the other baselines across all scenarios, offering a valuable and cost-effective solution for practical applications.
Ablation Studies.
In the ablation studies, we compare the performance of CAML with two other baselines: the fully-equipped teacher model, and a multi-agent BC model trained and tested using only RGB images, without knowledge distillation (KD). As shown in Table 2, while CAML shows a slight performance drop compared to the teacher, due to missing modalities at inference, it still achieves strong and robust performance. This highlights the effectiveness of our method in handling modality reduction scenarios, making it especially suitable for resource-constrained environments. Furthermore, CAML consistently outperforms the RGB-only BC baseline. The results demonstrate the advantage of incorporating multi-modal learning into multi-agent systems. Specifically, leveraging additional modalities during training through knowledge distillation allows CAML to encode richer information into the student model, resulting in improved performance at test time.
Table 2: Performance comparison (%) of CAML with the teacher model and the RGB-only BC baseline. While CAML shows a slight performance drop compared to the teacher model due to missing modalities at inference, it still achieves strong and robust performance. Moreover, CAML outperforms the RGB-only BC baseline across all scenarios.
Approach Overtaking Left Turn Red Light Violation
[…]
Teacher 96.0$\pm$0.3 86.8$\pm$0.4 73.3$\pm$0.8 81.2$\pm$0.6 68.8$\pm$0.9 85.8$\pm$0.7
RGB-only BC 85.2$\pm$1.0 83.2$\pm$0.9 56.4$\pm$1.5 78.0$\pm$1.2 55.8$\pm$1.3 81.0$\pm$1.2
CAML 92.8$\pm$0.8 85.5$\pm$0.5 67.3$\pm$1.0 79.7$\pm$1.1 66.1$\pm$1.4 83.0$\pm$1.3
@@
Overall, the experimental results clearly illustrate the superiority of our CAML framework. The ability of CAML to learn a more effective driving policy stems from the collaborative multi-modal behavior of multiple agents, which together capture a wider and more nuanced representation of data. This broader data coverage enables the ego vehicle to make better-informed decisions, improving safety and performance, particularly in complex, dynamic, and accident-prone environments where isolated agents with limited sensing.